# Pretraining and Updating Language- and Domain-specific Large Language Model: A Case Study in Japanese Business Domain

### v2

@@ 3.2 Compared Models and Instruction-tuning
We compare our pretrained model with six Japanese LLMs and three multilingual LLMs. Japanese LLMs are categorized into two: models pretrained from scratch and models with continual pretraining from a multilingual model. The former, that is from scratch, is the same setting as ours, and three out of the six models are used for the comparison; 1) llm-jp-13b [^9] https://huggingface.co/llm-jp/llm-jp-13b-v1.0 is pretrained on the Japanese Wikipedia, mC4 and English Wikipedia, The Pile (Gao et al., 2020), and The Stack (Kocetkov et al., 2022), 2) plamo-13b [^10] https://huggingface.co/pfnet/plamo-13b is pretrained on English texts from RedPajama and Japanese texts from mC4 and the Japanese Wikipedia, and 3) weblab-10b [^11] https://huggingface.co/matsuo-lab/weblab-10b is pretrained on Japanese texts from mC4 and English texts from The Pile. The first two are models with 13-billion parameters and the last has 10 billion parameters.
@@
We also compare our models trained by continual pretraining from existing multilingual models; 4) nekomata-14b [^12] https://huggingface.co/rinna/nekomata-14b is trained from qwen-14b on the Japanese CC100, C4, Wikipedia, Oscar, and Rinna curated corpus, and The Pile, 5) ELYZA-japanese-13b [^13] https://huggingface.co/elyza/ELYZA-japanese-Llama-2-13b is trained from Llama-2-13b on Japanese texts from Oscar and Wikipedia, 6) Swallow-13b [^14] https://huggingface.co/tokyotech-llm/Swallow-13b-hf is trained from Llama-2-13b on Japanese texts from Wikipedia and meticulously cleaned Common Crawl and English texts from RefinedWeb (Penedo et al., 2023) and The Pile. The model 4) has 14 billion while others have 13 billion parameters.
@@
We also use multilingual LLMs for comparison; for 7) gpt-3.5-turbo-0125, 8) gpt-4-1106-preview, models are selected from the OpenAI API. 9) Llama-2-13b-hf (Touvron et al., 2023) [^15] https://huggingface.co/elyza/ELYZA-japanese-Llama-2-13b is a 13-billion English LLM, which can produce Japanese texts.
LLMs without instruction tuning rarely provide accurate answers to these questions, which suggests the necessity of this tuning. There are multiple data choices for instruction tuning in Japanese, such as Ichikara, Alpaca, Dolly, and Oast. The statistics for these datasets are presented in Table 3. For the dataset selection, we conduct preliminary experiments using our proposed model detailed in Table 4. The values in the table demonstrate that all instruction-tuned models generally perform similarly, with average scores of approximately 0.550, except for the model trained on the Oast dataset, which achieves only 0.510. However, the model tuned on the Ichikara dataset consistently performs well on question-answering tasks, achieving scores of 0.78 on JSQuAD, 0.86 on JAQKET, and 0.84 on JCommonsenseQA. Therefore, we opt for the Ichikara dataset for instruction tuning.
Note that both our LLMs and the compared models are instruction-tuned on Ichikara dataset  [^16] https://liat-aip.sakura.ne.jp/wp/llmのための日本語インストラクションデータ作成/llmのための日本語インストラクションデータ-公開/. We use Low-Rank Adaptation (LoRA) Hu et al. (2021) for the instruction tuning.
@@
Oast 0.77 0.85 0.81 0.74 0.25 0.07 0.08 0.510
Table 4: Preliminary experiments: the comparison between datasets for instruction tuning. All base-model is our proposed 13-billion Japanese business domain specific LLM. JCom. and XW. refer to the JCommonSense and XWinograd datasets, respectively. The Ave. column shows the averaged values over all datasets.
@@ 4 Results
Results on Domain-specific Benchmark: Table 5 presents the results on the domain-specific benchmark. The table is divided into three categories representing different models: Japanese LLMs pretrained from scratch, Japanese LLMs with continual pretraining, and multilingual LLMs.
In the case of NoContext-QA, our pretrained model achieves the highest accuracy at $0.90$, compared to other existing Japanese LLMs such as nekomata-14b and Swallow-13b, which score $0.78$ for both. In this setting, the LLMs must generate answers from internally stored knowledge without using external knowledge such as retrieved documents. This suggests that our model possesses domain-specific knowledge within LLM.
For ManualRAG-QA, our model achieves the best score among the LLMs pretrained from scratch. For example, our model achieves an accuracy of $0.84$, whereas the other full-scratch models achieve accuracies of $0.76$, $0.82$, and $0.84$, respectively. Similar tendencies are observed in the AutoRAG-QA task. Our model achieves $0.74$ whereas the other models achieve only values of $0.62$, $0.64$, and $0.34$, respectively. These RAG tasks involve comprehension and rely on language ability to find answers in documents. This is in constrast to NoContext-QA, which tests the model’s ability to generate answers without document retrieval. From these results, we conclude that our pretrained model excels in answering business-specific questions because of its knowledge; however, the Japanese LLMs with continual training can achieve better performance in language comprehension tasks.
Comparing the two strategies for pretraining (from scratch or continual), continual pertinence consistently outperforms the models pretrained from scratch. Among the general-domain Japanese LLMs, the best performing LLMs with continual pretraining i.e., Swallow-13b and nekomata-14b, achieve $0.78$ in NoContext-QA, whereas the best general-domain LLM pretrained from scratch i.e., llm-jp-13b, achieves a notably lower score of $0.34$. Despite the fact that the full scratch strategy seems disadvantageous overall, it should be noted that our model performs the best on the NoContext-QA task among all compared models, which supports the effectiveness of the use of domain-specific data in pretraining. For the QAs with contexts i.e., ManualRAG-QA and AutoRAG-QA, our model achieves the best score ($0.84$ and $0.74$) among the full scratch models but the best performing models with continual pretraining perform better ($0.94$ and $0.82$). One of possible reasons is that continual models e.g., Swallow-13b or ELYZA-japanese-13b, use Llama2 as a base model, which is pretrained on a larger dataset. Thus, the base model might provide a strong language comprehension ability. Although this paper focuses on pretraining models from scratch, continually learning a strong existing model such as Llama2 on our domain-specific corpus may produce much more performance. We leave further investigations for future work.
@@
Results on General-domain Benchmark: Table 6 presents the results on the general-domain benchmark, “lm-evaluation-harness.” Among the models pretrained from scratch, our model demonstrates superior performance in terms of the averaged score i.e., ours achieves $0.55$ on average while the best performing Japanese LLMs pretrained from scratch achieve $0.49$. In particular, in QA tasks without context, such as JAQKET and JCommonSense, our model achieves the best scores $0.78$ and $0.85$, respectively.
@@
Table 6: Results on lm-evaluation-harness. JCom. and XW. refer to the JCommonSense and XWinograd datasets, respectively. The Ave. column shows the averaged values over all datasets. OpenAI’s GPTs cannot be used for this experiments because the model parameters are not public.
Analysis of Models with Latest Business Documents: We compare the updated model with the original pretrained one. The left part of Table 7 presents the validation loss values on two datasets: the 8,192 documents used for the validation portion in pretraining (Pretrain Data) and the latest business documents used for updating the model (Latest Data). The loss value of the updated model with $r=0.0$, indicating it is trained only on the latest documents, is 2.05 on the latest data. This is an improvement over the original pretrained model’s $2.25$. This indicates that the updated model better fits the latest documents. However, the loss on the pretaining data worsens from 2.11 to 2.19, compared to the original model. Increasing the value of $r$, that is adding more pretraining data to the latest documents, can alleviate the increase in the loss of the pretraining data. When we increased $r$ to $0.3$, we achieve a loss of 2.12, nearly equivalent to the original model’s 2.11, without degrading the loss on the latest documents (2.04).
To investigate the capability of our model to answer questions that require the latest knowledge. We create a set of 10 questions about latest business topics, specifically from October to November 2023. We selected target topics that showed a notable increases in access to a search engine, compared to that in September 2023. The accuracy of NoContext-QA on the non-latest questions and latest questions is shown in the right part of Table 7. In the latest questions, the updated models demonstrate higher score ($0.90$) than the pretrained model ($0.30$), denoting that a Japanese- and business-specific model can learn new information with the continual learning. Different from the results in terms of loss, we did not observe significant changes in terms of accuracy among the models with different values of $r$.
@@
Examples of Outputs Table 8 shows the examples of the outputs from our pretrained model and compared LLMs. When we input a question “Which two banks failed in March 2023?”, our pretrained model correctly generates two banks: the Silicon Valley Bank and the Signature Bank. Other three Japanese LLMs’ outputs contain hallucinations, which are indicated in bold in the table. For example, Wells Forgo, Bank of America, and Citigroup did not fail in March 2023. We observed a similar tendencies for the other questions.
@@
Ours The two U.S. banks that failed in March 2023 were Silicon Valley Bank and Signature Bank. (2023年3月に経営破綻した米国の2 つの銀行は、シリコンバレー銀行とシグネチャー銀行である。)
llm-jp-13b-v1.0 The two U.S. banks that failed in March 2023 were Wells Fargo and Bank of America. (2023年3月に経営破綻したアメリカの2つの銀行は、ウェルズ・ファーゴとバンク・オブ・アメリカである。
plamo-13b The two banks that failed in March 2023 were Bank of America and Citigroup. (2023年3月に経営破綻した2つの銀行は、バンク・オブ・アメリカとシティ・グループでした。)
weblab-10b The two U.S. banks that failed in March 2023 were Bank of America and Wells Fargo. (2023年3月に経営破綻したアメリカの2つの銀行は、Bank of America と Wells Fargoです。)
ChatGPT (gpt-3.5-turbo-0125) The two American banks that went bankrupt in March 2023 are Regional Asahi Bank and Centry National Bank. (2023年3月に経営破綻したアメリカの2つの銀行は、リージョナル朝陽銀行とセントリーナショナル銀行です。)
Table 8: Examples of output generated from different LLMs. The sentences in the buckets are the original outputs in Japanese. Hallucinated terms are presented in bold.
@@ 5 Related Work
The language-specific datasets for instruction tuning have been released. For Japanese, Ichikara, Alpaca, Dolly, and Oast are common; however, the performances of models trained on these datasets have not been compared in depth. Our preliminary experiment is the first to compare these datasets, which can be considered as an important contribution for the Japanese LLM community. Also, our experiments can provide insights for researchers who focus on other non-English languages.
Domain-specific models are pretrained on two different ways; training from scratch or continual training. Representative examples of the former started for the encoder-only models such as SciBERT (Beltagy et al., 2019) in the science domain and later decoder-only large language models follow, e.g., BloombergGPT (Wu et al., 2023). The continually learned domain-specific LMs include BioBERT (Lee et al., 2019), exBERT (Wang et al., 2020). Scratch approaches are often been used for settings where we can get sufficient data. Our setting is categorized into this setting, thus, we use the scratch approach.
The combination of the two directions, i.e., language- and domain-specific settings, is underexplored in Japanese. Japanese domain-specific language model does not exist, except for the pretrained BERT for the material science domain (Ishigaki et al., 2023). Our target is decoder-only architecture with many more parameters for the business domain, which has high demand in the industry but is less studied. Our pretrained model is the first Llama-based domain-specific model for Japanese
Continual pretraining is a promising direction if we have only a small dataset for pretraining. Our experiments for updating the pretrained model are categorized in this setting. “Catastrophic forgetting” is a major problem in this case (Ling et al., 2023). Scialom et al. (2022) suggest that the mixture of two types of data can mitigate this problem, thus, we used the technique to mix the latest documents and older ones.
@@ 6 Conclusion
This paper presented the first Japanese business domain-specific LLM. We pretrain the LLM with 13 billion parameters from scratch and will be made public. Evaluations demonstrated that our LLM achieves better on the domain-specific benchmark we newly released. We also found that updating the pretrained parameters with the latest business documents can make the model better fit the latest documents. For future work, we will explore methods that leverage the context in RAG setting while utilizing the knowledge within LLMs.
@@ A.1 Templates used in Experiments
Table 9: Examples of output generated from different LLMs. We show the English translations while the sentences in the buckets are the original outputs in Japanese. Hallucinated terms are in bold.
@@ A.2 Prompts used in Experiments
Please answer to the given question. If the answer to the question is included in the article text, please use the answer from the text. If the article does not contain the answer please state that ”the article does not contain the answer” and answer the question using your knowledge. Question:{question} Article Text: {the first 1k characters} ### Output:

### v3

@@ 3.2 Compared Models and Instruction Tuning
Among the models pretrained from scratch, we compare three: 1) llm-jp-13b(LLM-jp et al., 2024), which is pretrained on Japanese Wikipedia, mC4, English Wikipedia, The Pile (Gao et al., 2020), and The Stack (Kocetkov et al., 2022); 2) plamo-13b [^9] https://huggingface.co/pfnet/plamo-13b, which is pretrained on English texts from RedPajama and Japanese texts from mC4 and Japanese Wikipedia; and 3) weblab-10b [^10] https://huggingface.co/matsuo-lab/weblab-10b, which is pretrained on Japanese texts from mC4 and English texts from The Pile. The first two models have 13 billion parameters, while the last has 10 billion.
@@
We also include models that were continually pretrained from existing multilingual models: 4) nekomata-14b(Sawada et al., 2024), which is trained from qwen-14b on Japanese CC100, C4, Wikipedia, Oscar, Rinna curated corpus, and The Pile; 5) ELYZA-japanese-13b [^11] https://huggingface.co/elyza/ELYZA-japanese-Llama-2-13b, which is trained from Llama-2-13b on Japanese texts from Oscar and Wikipedia; and 6) Swallow-13b(Fujii et al., 2024), which is trained from Llama-2-13b on Japanese texts from Wikipedia, meticulously cleaned Common Crawl, and English texts from RefinedWeb (Penedo et al., 2023) and The Pile. Models 4) has 14 billion parameters, while the others have 13 billion.
@@
Additionally, we compare our models with multilingual LLMs: 7) gpt-3.5-turbo-0125, 8) gpt-4-1106-preview, selected from the OpenAI API; and 9) Llama-2-13b-hf (Touvron et al., 2023) [^12] https://huggingface.co/elyza/ELYZA-japanese-Llama-2-13b, a 13-billion parameter English LLM capable of generating Japanese text.
LLMs that are not instruction-tuned often struggle to provide accurate answers, highlighting the importance of instruction-tuning. We evaluated several datasets for instruction-tuning in Japanese, including Ichikara (Sekine et al., 2023), Alpaca (Taori et al., 2023; Shimizu, 2023), Dolly (Conover et al., 2023; Kuniyoshi, 2023a), and OASST (Köpf et al., 2023; Kuniyoshi, 2023b). Statistics for these datasets are provided in Table 3. Preliminary experiments shown in Table 4 reveal that the use of Ichikara dataset consistently performs well on question-answering tasks, with scores of 0.78 on JSQuAD, 0.86 on JAQKET, and 0.84 on JCommonsenseQA. Therefore, we chose the Ichikara dataset for instruction-tuning our LLMs. We use Low-Rank Adaptation (LoRA) (Hu et al., 2021) for the instruction tuning.
@@
OASST 0.77 0.85 0.81 0.74 0.25 0.07 0.08 0.510
Table 4: Preliminary experiments: the comparison between datasets for instruction tuning. All base-model is our proposed 13-billion Japanese business domain specific LLM. JCom. and XW. refer to the JCommonsenseQA and XWinograd datasets, respectively. The Ave. column shows the averaged values over all datasets.
@@ 4 Results
4.1 Results on Business Question Benchmark:
Table 5 displays the results from our domain-specific benchmark, categorizing the models into three groups: Japanese LLMs pretrained from scratch, Japanese LLMs with continual pretraining, and multilingual LLMs.
In the NoContext-QA setting, our pretrained model achieves the highest accuracy of $0.90$. This is notably higher compared to other Japanese LLMs such as nekomata-14b and Swallow-13b, which both score $0.78$. This performance indicates that our model effectively utilizes domain-specific knowledge stored internally without relying on external documents.
For the ManualRAG-QA task, our model again performs best among LLMs pretrained from scratch, achieving an accuracy of $0.84$. In contrast, the other models in this category score $0.76$ (llm-jp-13b), $0.82$ (plamo-13b), and $0.52$ (weblab-10b). A similar trend is observed in the AutoRAG-QA task, where our model scores $0.74$, while other models score $0.62$ (llm-jp-13b), $0.64$ (plamo-13b), and $0.34$ (weblab-10b), respectively. In these RAG settings, which require comprehension and the ability to extract information from documents, our models perform better than other full-scratch LLMs. However, Japanese LLMs with continual pretraining, such as ELYZA-japanese-13b and Swallow-13b, achieve higher scores of $0.94$ and $0.90$ in ManualRAG-QA, respectively. These values suggest that continual pretraining works better in comprehension.
When comparing pretraining strategies, continual pretraining outperforms full-scratch models except for our model. For instance, in the NoContext-QA task, models with continual pretraining, such as Swallow-13b and nekomata-14b, achieve scores of $0.78$, whereas the best pretrained full-scratch baseline, llm-jp-13b, scores $0.34$. Despite the overall advantage of continual pretraining, our model performs the best in the NoContext-QA task, highlighting the value of domain-specific data to strengthen the internal domain-specific knowledge in LLMs.
For tasks involving context, such as ManualRAG-QA and AutoRAG-QA, our model achieves the highest scores of $0.84$ and $0.74$, respectively, among full-scratch models. However, models with continual pretraining outperform in these tasks with scores of $0.94$ and $0.82$. This may be attributed to the use of Llama2 as a base model for continual pretraining, which benefits from being pretrained on a larger dataset and thus has stronger language comprehension capabilities.
Although this study focuses on models pretrained from scratch, future work could explore the potential performance gains from continually training a strong existing model, such as Llama2, on a domain-specific corpus.
@@ 4.1 Results on Business Question Benchmark:
4.2 Results on General-domain Benchmark:
Table 6 presents the results on the general-domain benchmark, “lm-evaluation-harness.” Among the models pretrained from scratch, our model demonstrates superior performance in terms of the averaged score i.e., ours achieves $0.55$ on average while the best performing Japanese LLMs pretrained from scratch achieve $0.49$. In particular, in QA tasks without context, such as JAQKET and JCommonsenseQA, our model achieves the best scores $0.78$ and $0.85$, respectively.
@@ 4.2 Results on General-domain Benchmark:
Table 6: Results on lm-evaluation-harness. JCom. and XW. refer to the JCommonsenseQA and XWinograd datasets, respectively. The Ave. column shows the averaged values over all datasets. OpenAI’s GPTs cannot be used for this experiments because the model parameters are not public.
4.3 Analysis of Updated Models
We compare the updated model with the original pretrained model. The left part of Table 7 shows the validation loss values on two datasets: the 8,192 documents that were used for the validation portion during pretraining (Pretrain Data) and the latest business documents used for updating the model (Latest Data). If our LLM were able to learn the latest knowledge without losing general knowledge, it would achieve low validation loss on both Pretrain Data and Latest Data. The loss value for the updated model with $r$ = 0.0, indicating that it was trained solely on the latest documents, is 2.05 on Latest Data. This represents an improvement over the original pretrained model’s loss of 2.25, suggesting that the updated model acquired the latest knowledge better. However, the loss on the Pretrain Data increased from 2.11 to 2.19 compared to the original model, which suggests the worse fit to the non-latest documents.
Increasing the value of $r$, which incorporates more pretraining data along with the latest documents, mitigates the increase in loss on the pretraining data. For example, with $r$ set to 0.3, the loss is 2.12 on Pretrain Data, which is nearly equivalent to the original model’s loss of 2.11, while the loss on the latest documents remains stable at 2.04.
The right part of Table 7 shows the accuracy of our compared updated LLMs in question answering about latest news. We created a set of 10 questions (LatestQ) about recent business topics from October to November 2023. We selected topics that had shown a notable increase in search engine access compared to September 2023. As result, all the updated models achieve a higher accuracy (0.90) compared to the pretrained model (0.30), indicating that a Japanese- and business-specific model can acquire new information through continual learning. Unlike the results regarding loss, we did not observe significant differences in accuracy among models with different values of $r$ on the Business Question Benchmark (Non-LatestQ) where we obtained high values e.g., $0.90$ (Pretrained model and Updated model($r=0.3$)) or $0.92$ (Updated model ($r=0.0,0.1$)). We conclude that using only the latest articles can lead to degradation in both loss and accuracy and incorporating around 10% of non-latest articles proves effective, while including 30% is excessive.
@@ 4.3 Analysis of Updated Models
4.4 Examples of Outputs
Table 8 shows the examples of the outputs from our pretrained model and compared LLMs. When we input a question “Which two banks failed in March 2023?”, our pretrained model correctly generates two banks: the Silicon Valley Bank and the Signature Bank. Other three Japanese LLMs’ outputs contain hallucinations, which are indicated in bold in the table. For example, Wells Forgo, Bank of America, and Citigroup did not fail in March 2023. We observed a similar tendency for the other questions.
Question : Which two American banks went bankrupt in March 2023?
@@ 4.4 Examples of Outputs
Ours The two U.S. banks that failed in March 2023 were Silicon Valley Bank and Signature Bank. ({CJK}UTF8ipxg2023年3月に経営破綻した米国の2つの銀行は、シリコンバレー銀行とシグネチャー銀行である。))
llm-jp-13b-v1.0 The two U.S. banks that failed in March 2023 were Wells Fargo and Bank of America. ({CJK}UTF8ipxg2023年3月に経営破綻したアメリカの2つの銀行は、ウェルズ・ファーゴとバンク・オブ・アメリカである。)
plamo-13b The two banks that failed in March 2023 were Bank of America and Citigroup. ({CJK}UTF8ipxg2023年3月に経営破綻した2つの銀行は、バンク・オブ・アメリカとシティ・グループでした。)
weblab-10b The two U.S. banks that failed in March 2023 were Bank of America and Wells Fargo. ({CJK}UTF8ipxg2023年3月に経営破綻したアメリカの2つの銀行は、Bank of America と Wells Fargoです。)
ChatGPT (gpt-3.5-turbo-0125) The two American banks that went bankrupt in March 2023 are Regional Asahi Bank and Centry National Bank. ({CJK}UTF8ipxg2023年3月に経営破綻したアメリカの2つの銀行は、リージョナル朝陽銀行とセントリーナショナル銀行です。)
Table 8: Translated examples of output generated from different LLMs. The sentences in the buckets are the original outputs in Japanese. Hallucinated terms are presented in bold.
@@ 5 Related Work
The language-specific datasets for instruction tuning have been released. For Japanese, Ichikara, Alpaca, Dolly, and OASST are common; however, the performances of models trained on these datasets have not been compared in depth. Our preliminary experiment is the first to compare these datasets, which can be considered as an important contribution for the Japanese LLM community. Also, our experiments can provide insights for researchers who focus on other non-English languages.
Domain-specific models are pretrained in two different ways; training from scratch or continual training. Representative examples of the former started for the encoder-only models such as SciBERT (Beltagy et al., 2019) in the science domain and later decoder-only large language models follow, e.g., BloombergGPT (Wu et al., 2023). The continually learned domain-specific LMs include BioBERT (Lee et al., 2019), exBERT (Wang et al., 2020). Scratch approaches are often used for settings where we can get sufficient data. Our setting is categorized into this setting, thus, we use the scratch approach.
The combination of the two directions, i.e., language- and domain-specific settings, is underexplored in Japanese. Japanese domain-specific language model does not exist, except for the pretrained BERT for the material science domain (Ishigaki et al., 2023). Our target is decoder-only architecture with many more parameters for the business domain, which has high demand in the industry but is less studied. Our pretrained model is the first Llama-based domain-specific model for Japanese.
Continual pretraining is a promising direction if we have only a small dataset for pretraining. Our experiments for updating the pretrained model with the latest news are categorized in this setting. “Catastrophic forgetting” is a major problem in this case (Ling et al., 2023). Scialom et al. (2022) suggest that the mixture of two types of data can mitigate this problem, thus, we used the technique to mix the latest documents and older ones.
@@ 6 Conclusion
This paper presented the first Japanese business domain-specific LLM. We pretrain the LLM with 13 billion parameters from scratch and the model is released to be publicly available. We also update the model parameters by the latest articles and confirmed performance gains. For updates, we conclude that using only the latest articles can lead to performance degradation but incorporating around 10% of non-latest articles proves effective. Comprehensive evaluations demonstrated that our LLM achieves the best accuracy score without retrieval on the domain-specific benchmark we newly released. These results provide valuable insights for researchers working on other domains and languages. For future work, we will explore comparing different ways to train language- and domain-specific LLMs e.g., full-stratch v.s. continual pretraining.
@@ A.1 Templates used in Experiments
Table 9: Settings of the template of the lm-evaluation-harness. We experimented every model in Table 6 with the templates listed here. The scores reported in Table 6 are the highest among the variants of the templates.
@@ A.2 Prompts used in Experiments
Please answer to the given question. If the answer to the question is included in the article text, please use the answer from the text. If the article does not contain the answer please state that "the article does not contain the answer" and answer the question using your knowledge. Question:{question} Article Text: {the first 1k characters} ### Output: